realman:

Select all that apply. Which steps are involved in multiplying fractions? Find a common denominator Multiply the numerators together Check to see if the product can be simplified Multiply the denominators together Convert the product to a mixed number

jhonyy9:

so for your help just an example \[\frac{ a }{ b }*\frac{ c }{ d} = \frac{ ac }{ bd}\] hope helped understandably